# Pushing the Limits of Sparsity: A Bag of Tricks for Extreme Pruning

### v2

@@ Abstract
Pruning of deep neural networks has been an effective technique for reducing model size while preserving most of the performance of dense networks, crucial for deploying models on memory and power-constrained devices. While recent sparse learning methods have shown promising performance up to moderate sparsity levels such as 95% and 98%, accuracy quickly deteriorates when pushing sparsities to extreme levels. Obtaining sparse networks at such extreme sparsity levels presents unique challenges, such as fragile gradient flow and heightened risk of layer collapse. In this work, we explore network performance beyond the commonly studied sparsities, and propose a collection of techniques that enable the continuous learning of networks without accuracy collapse even at extreme sparsities, including 99.90%, 99.95% and 99.99% on ResNet architectures. Our approach combines 1) Dynamic ReLU phasing, where DyReLU initially allows for richer parameter exploration before being gradually replaced by standard ReLU, 2) weight sharing which reuses parameters within a residual layer while maintaining the same number of learnable parameters, and 3) cyclic sparsity, where both sparsity levels and sparsity patterns evolve dynamically throughout training to better encourage parameter exploration. We evaluate our method, which we term Extreme Adaptive Sparse Training (EAST) at extreme sparsities using ResNet-34 and ResNet-50 on CIFAR-10, CIFAR-100, and ImageNet, achieving significant performance improvements over state-of-the-art methods we compared with.
@@ 1 Introduction
Network pruning [9, 10, 16, 22, 19, 15] is a widely-used technique for reducing a network’s parameters and compressing its size. Reducing model sizes is crucial for deploying models on edge devices with limited resources. Conventionally, pruning methods have focused on reducing parameters from pre-trained models. However, it requires at least as much computation as training a dense model as it must converge before pruning takes place. The Lottery Ticket Hypothesis [6, 7, 23] has shown that there exist subnetworks that can perform as well as the dense counterpart, even when trained from an initially sparse state. This insight has recently gained much traction in sparse training, a paradigm in which sparse networks are trained from scratch without the need for dense pre-training, promising efficiency gains in both memory and computation.
Sparse training methods can be broadly classified into two categories: static sparse training (SST), also sometimes referred to as Pruning at Initialization (PAI), where the sparsity pattern is pre-determined at initialization and remains fixed throughout training [17, 33, 30, 2], and dynamic sparse training (DST), where the sparsity pattern continuously evolves during training [25, 3, 4, 34]. They have shown promising results at retaining most of the performance at moderate sparsity levels, commonly up to 98% in recent literature. However, most of these begin experiencing rapid degradation in accuracy beyond these levels due to limitations such as layer collapse and rapid gradient degradation [31, 33]. While the focus has largely been on preserving performance at moderate sparsities, there is an increasing demand to push towards ’extreme sparsities’, where achieving substantial model compression at the cost of performance trade-offs may be necessary. This shift is particularly crucial in resource-constrained settings, where aggressive sparsification can have substantial computational and memory benefits, even if some accuracy must be sacrificed. Despite this need, extreme sparsities remain relatively understudied, with only a few methods attempting to maintain usable accuracy at these extreme levels [31, 2, 28].
@@
  • We perform extensive empirical evaluation of EAST on computer vision tasks, and demonstrate that it can achieve superior performance at sparsities beyond 99.90%, surpassing all previous methods tested.
@@ 2.1 Dynamic Sparse Training
Early work by Mocanu et al. demonstrated the viability of DST through SET [25], which used magnitude pruning and random weight regrowth. Subsequent criteria for regrowing connections have been developed, such as using the momentum of parameters in SNFS [3] or their gradient in RigL [4] to determine the salience of deactivated connections. GraNet [20] adapted RigL’s approach by starting with half the parameters and gradually increasing to target sparsity instead of starting from the target sparsity. ITOP [21] provides a deeper understanding of pre-existing DST methods and enhances parameter exploration during sparse training. NeurRev [18] addressed the issue of dormant neurons by removing harmful negative weights. Recent advances include Top-KAST [13], which maintains an auxiliary set of weights for enhanced exploration, and BiDST [14] which formulates DST as a bi-level optimization problem for joint weight-mask optimization. MEST [36] introduces a memory-efficient sparse training framework for edge devices, while Chase [35] adapts dynamic unstructured sparsity into hardware-friendly channel-wise sparsity, both optimizing performance on resource-constrained systems without sacrificing accuracy. BiDST [14] formulates DST as a bi-level optimization problem, where both weights and masks are optimized simultaneously resulting in higher mask exploration efficiency. Supticket [34] further enhances the performance of DST by leveraging weight averaging at the late training stage.
2.2 Extreme Sparsity Methods
@@
Most current methods focus on sparsity levels at which pruned networks can match or closely approximate the performance of their dense counterparts [8]. However, there is growing interest in extreme sparsities, where accuracy trade-offs become inevitable but potentially worthwhile for substantial computational gains. Yet there remains challenges, such as maintaining stable gradient flow and preventing layer collapse, which conventional pruning methods struggle to address. GraSP [33] and SynFlow [31] specifically addressed the gradient preservation and layer collapse issues at high sparsities via careful network initialization. SNIP [17] and SNIP-it [32] proposed PaI methods that showed promising results at higher sparsities, though primarily focused on static masks. De Jorge et al. [2] demonstrated that methods like SNIP and GraSP can perform worse than random pruning at 99% sparsity and beyond, and proposed iterative pruning approaches (Iter-SNIP and FORCE) that maintained meaningful performance up to 99.5% sparsity through gradual parameter elimination. DCTpS [28] introduced ”DCT plus Sparse” layers that combined a fixed dense DCT offset matrix with trainable sparse parameters to maintain propagation at extreme sparsities up to 99.99%.
2.3 Activation Functions
@@
Activation functions in sparse networks have traditionally relied on static activation functions such as ReLU [27], which applies a fixed threshold to the input. Parametric variants such as PReLU [11] introduced learnable parameters to adapt the negative slope, offering slight improvements in gradient flow. Swish [29] and Mish [24] aim to smooth gradients for better optimization but remain static throughout training. Dynamic ReLU (DyReLU) [1] introduced an adaptive activation that dynamically adjusts the slopes of a ReLU based on the input.
@@ 3.1 Activation Function
In this work, when initializing a network, we replace the standard ReLU activation with Dynamic ReLU (DyReLU), specifically the DyReLU-B variant, which dynamically adapts and adjusts channel-wise activation coefficients during training, allowing the model to explore a larger space of activation functions. Preliminary details about DyReLU and DST related to this section can be found in Supplementary Material. At extreme sparsities (e.g., $>$ 99%), networks rapidly start losing expressive power. DyReLU compensates for this by dynamically adjusting the activation functions to adapt to the reduced network capacity. By using this during the early stages of training, we allow the network to explore a larger activation space and prevent premature collapse due to over-pruning.
We hypothesize that DyReLU can help networks better explore and optimize parameters at extreme sparsities by dynamically adjusting the activations during the critical early phases of training. This allows the network to better handle the challenges introduced by the sparse structure (e.g., loss of representational capacity), as the dynamic activations can help identify and amplify important features in the sparse subnetwork. We detail the ablation studies on the effect of DyReLU in section 4.3.
Phasing out DyReLU:
$\beta(t)=1-\frac{t-t_{\text{start}}}{t_{\text{end}}-t_{\text{start}}}$
$f(x;t=0)=f_{DyReLU}(x)$
$f(x;t)=\beta(t)\cdot f_{DyReLU}(x)+(1-\beta(t))\cdot f_{ReLU}(x)$
By the end of phasing $t=T_{\text{end}}$, we have:
$f(x;t=T_{\text{end}})=f_{ReLU}(x)$
Theorem:
During DyReLU to ReLU transition with decay factor $\beta(t)$, the gradient flow is strictly greater than standard ReLU such that:
$\|\nabla\mathcal{L}\|_{\text{DyReLU}\to\text{ReLU}}>\|\nabla\mathcal{L}\|_{\text{ReLU}}\quad\text{for }\beta(t)\geq 0$
The proof for this can be found in Supplementary Materials. In the experiments, we show how this method initially boosts the gradient flow, and then keeps it even after $\beta$ reaches zero.
@@ 3.2 Weight Sharing
Much research in sparse training has been about path optimization - that is obtaining the right mask, or the right set of learnable parameters that are most important to the outcome. We ask the question: Can increasing the number of paths for gradient propagation while maintaining number of learnable parameters help maintain stable training at extreme sparsities? We hypothesize that creating more paths for gradient propagation through shared parameters can strengthen gradient flow, and lead to stronger overall gradient flow.
To test our hypothesis, we introduce a weight sharing mechanism across residual blocks within each layer that preserves gradient flow without increasing the overall count of learnable parameters. Specifically, we choose a block inside a layer and share its parameters (along with their masks) with all blocks after it. These blocks reference the single block tensor that they share with, and multiply by a scaling factor for forward and backward passes. They do not need to be physically stored multiple times in memory, making the approach memory-efficient. Thus, we adjust sparsity calculation to reflect this architectural modification, such that sparsity $\text{sparsity s}=1-\frac{\lVert\mathcal{M}\odot\theta_{\text{learnable}}\rVert_{0}}{\lVert\theta\rVert_{0}}$, where $\lVert$.$\rVert$ . ∥ is the $\ell_{0}$-norm or the cardinality, $\theta_{\text{learnable}}$ is the learnable parameters, and $\theta$ is the theoretical parameters in the original network with a mask $\mathcal{M}\in\{0,1\}^{|\theta|}$. In our experiments, we share the second block with the rest of the blocks within a layer. For instance, as shown in Figure 3, if a layer has 4 blocks, the second block shares its parameters and sparse topology with the 3rd and 4th block.
Figure 3: A layer in ResNet-34 with 4 blocks. Block 3 and Block 4 share parameters (and the masks) with block 2. Specifically, conv1 layer (green) of block 3 reuses conv1 layer of block 2, and is multipled by a learnable scaling factor; similarly, its conv2 layer (blue) reuses conv2 layer of block 2, and is multiplied by another scaling factor.
Formally, let $L$ be the number of residual blocks in a layer, and $R$ be the block that shares parameters ($1<R\leq L$), where the remaining $L-R$ blocks share parameters with the $R$-th block. Let $\theta_{i}$ represent the parameters of the $i$-th block. Our weight sharing scheme can be expressed as:
@@ 3.3 Cyclic Sparsity
While standard DST methods such as SET and RigL adjust the sparsity pattern, sparsity $s$ is consistent throughout training - pruning and regrowing of weights take place at the same $t$. i.e., $\mathcal{M}(t)=\text{Prune}(\theta(t))\cup\text{Grow}(\nabla_{\theta}\mathcal{L})$.
@@
In our work, we introduce a cyclic sparsity schedule where the network’s sparsity level itself evolves dynamically during training. Instead of maintaining a constant sparsity, we cyclically adjust the sparsity $s(t)$ between a maximum value $s_{\text{max}}$ and a minimum value $s_{\text{min}}$. The sparsity level $s(t)$ is governed by a cosine schedule:
@@ 4 Experiments
Our experiments include image classification using ResNet34 and ResNet50 [12] on computer vision datasets CIFAR-10 and CIFAR-100, and ResNet50 on ImageNet. We compare our method with DST methods SET[25] and RigL[4], and a competitive SST method SynFlow[31], which also experiments at very high sparsity. We collectively refer to these methods as sparse training methods for the rest of the paper. For SynFlow, we use the official repository with its default hyperparameters to evaluate its performance. For SET and RigL, we use the repository provided by ITOP [21]. We keep all hyperparameters as they are optimally configured in the paper, including $\Delta T$, ERK initialization, the mask update interval of 1500 and 4000 for SET and RigL, respectively.
@@
4.1 CIFAR-10 and CIFAR-100
We evaluate EAST on CIFAR-10 and CIFAR-100 classification tasks using ResNet34 and ResNet50, with a particular focus on extreme sparsities (99%-99.99%). We repeat our experiments three times and report the average accuracy in Table 1, which includes the comprehensive results across different configurations.
At 99%, all dynamic training methods achieve comparable performance, significantly outperforming SynFlow. This is unsurprising as a static mask is expected to lead to worse performance than a dynamic mask [4, 26]. EAST’s advantages become increasingly pronounced at higher sparsities. At 99.90% and 99.95%, EAST beats the next best performing method RigL by a clear margin ranging between 1-2% with ResNet34 on both datasets, showing significant gains in performance. The difference becomes even more clear with ResNet50, where other methods take a huge plummet on accuracy and EAST shows a more gradual decline. EAST displays an improvement of 38-40% over RigL on CIFAR-10, and 1-7% on CIFAR-100 at 99.90% and 99.95%.
The most striking difference appears at 99.99%. We observe that all other methods result in complete collapse to random performance on either dataset and model. In contrast, EAST maintains meaningful accuracies with the exception of ResNet50 on CIFAR-100. For instance, EAST achieves 70.57% on CIFAR-10 compared to 10% random baselines. This pattern holds across both models, though ResNet50 shows more degradation at the highest sparsity levels, presumably due to more prominent vanishing gradient problems in a deeper network, with even fewer parameters spread across each layer.
We evaluate DCT and find that while DCT performs significantly better at 99.95% and 99.99%, it performs worse than SET, RigL and EAST at 99% and sometimes 99.9%. For instance, EAST achieves 3-8% higher accuracy than DCT at 99%, without involving the additional overhead. We further combine DCT and EAST and observe that EAST provides consistent improvements across all configurations tested. This demonstrates its versatility and complementary benefits to specialized extreme sparsity methods.
4.2 ImageNet
Method Top-1 Acc Top-5 Acc Top-1 Acc Top-5 Acc Throughput Latency
(%) (%) (%) (%) (imgs/sec) (ms)
Sparsity 99.50% Sparsity 99.90% Inference
@@ 4.2 ImageNet
DCT+RigL 19.99 43.34 6.62 18.84 1229.25 10.18
EAST (ours) 34.45 62.44 11.23 29.47 1638.41 6.80
@@
Table 2: Performance comparison on ImageNet at extreme sparsity levels using ResNet50. For throughput and latency measurements, we use batch sizes of 2 and run 100 times and report the average.
@@
Prior methods like SynFlow and DCTpS were evaluated primarily on smaller datasets like CIFAR and Tiny-ImageNet. We extend these baselines to ImageNet through implementation and adaptation of their published code and evaluate the scalability of EAST on ImageNet using ResNet50 at sparsity 99.5% and 99.9%. Given the much shortened training epochs for ImageNet, we implement only the DyReLU phasing and weight sharing components, excluding cyclic sparsity. As presented in Table 2, EAST has demonstrated capability in maintaining network robustness on large-scale dataset, outperforming previous SOTA by a significant margin, achieving approximately 2% and 2.7% higher top-1 accuracy than RigL. This advantage is more notable when comparing to other methods including SynFlow and SET, which fail to maintain meaningful performance and collapase to random accuracy.
@@
Interestingly, we observe different dynamics on ImageNet compared to CIFAR experiments. While DCTpS showed competitive advantages on smaller datasets, it does not exhibit the same competitive advantages as seen before, even with the extra overhead. DCTpS alone results in performance collapse and it performs worse than RigL. When combined with RigL, the accuracy improvements are modest and may be attributed primarily to RigL’s robust dynamic sparse exploration policy, rather than the DCT offset. On the inference efficiency front, EAST demonstrates practical advantages. Unlike DCTpS, which relies on a dense DCT matrix, the model resulted from EAST does not have such drawback, as all DyReLU activations have been replaced with regular ReLU and thus runs just like the original ResNet model with ReLU. Our inference measurements show 1.5X lower latency and 1.33X higher throughput compared to DCTpS, while simultaneously achieving better accuracy. These results demonstrate that EAST can scale up to large datasets while maintaining computational efficiency and accuracy advantages over existing methods.

### v3

@@ Abstract
Pruning of deep neural networks has been an effective technique for reducing model size while preserving most of the performance of dense networks, crucial for deploying models on memory and power-constrained devices. While recent sparse learning methods have shown promising performance up to moderate sparsity levels such as 95% and 98%, accuracy quickly deteriorates when pushing sparsities to extreme levels due to unique challenges such as fragile gradient flow. In this work, we explore network performance beyond the commonly studied sparsities, and propose a collection of techniques that enable the continuous learning of networks without accuracy collapse even at extreme sparsities, including 99.90%, 99.95% and 99.99% on ResNet architectures. Our approach combines 1) Dynamic ReLU phasing, where DyReLU initially allows for richer parameter exploration before being gradually replaced by standard ReLU, 2) weight sharing which reuses parameters within a residual layer while maintaining the same number of learnable parameters, and 3) cyclic sparsity, where both sparsity levels and sparsity patterns evolve dynamically throughout training to better encourage parameter exploration. We evaluate our method, which we term Extreme Adaptive Sparse Training (EAST) at extreme sparsities using ResNet-34 and ResNet-50 on CIFAR-10, CIFAR-100, and ImageNet, achieving significant performance improvements over state-of-the-art methods we compared with.
@@ 1 Introduction
Network pruning [9, 10, 16, 22, 19, 15] is a widely-used technique for reducing a network’s parameters and compressing its size. Reducing model sizes is crucial for deploying models on edge devices with limited resources. Conventionally, pruning methods have focused on reducing parameters from pre-trained models. However, it requires at least as much computation as training a dense model as it must converge before pruning takes place. The Lottery Ticket Hypothesis [6, 7, 23] gives theoretical foundation that subnetworks have the potential to reach full performance even when trained from an initially sparse state. This insight has recently gained much traction in sparse training, a paradigm in which sparse networks are trained from scratch without the need for dense pre-training.
Sparse training methods can be broadly classified into two categories: static sparse training (SST), also sometimes referred to as Pruning at Initialization (PaI), where the sparsity pattern is pre-determined at initialization and remains fixed throughout training [17, 33, 30, 2], and dynamic sparse training (DST), where the sparsity pattern continuously evolves during training [25, 3, 4, 34]. These methods perform well at moderate sparsity commonly up to 98%, but experience rapid accuracy degradation beyond this due to limitations such as layer collapse and gradient issues [31, 33]. There is an increasing demand to push towards ’extreme sparsities’, particularly in resource-constrained settings where substantial computational and memory benefits may justify some accuracy trade-offs. Our work is motivated by this practical consideration, and thus we emphasize that our primary focus is on the deployable model used for inference rather than training efficiency. Despite this need, extreme sparsities remain relatively understudied, with only a few methods attempting to maintain usable accuracy at these extreme levels [31, 2, 28].
@@
  • We perform extensive empirical evaluation of EAST on image classification tasks, and demonstrate that it can achieve superior performance at sparsities beyond 99.90%, surpassing all previous methods tested.
@@ 2 Related Work
Dynamic Sparse Training. Early work by Mocanu et al. demonstrated the viability of DST through SET [25], which used magnitude pruning and random weight regrowth. Subsequent criteria for regrowing connections have been developed, such as using the momentum of parameters in SNFS [3] or their gradient in RigL [4] to determine the salience of deactivated connections. GraNet [20] adapted RigL’s approach by starting with half the parameters and gradually increasing to target sparsity instead of starting from the target sparsity. ITOP [21] provides a deeper understanding of pre-existing DST methods and enhances parameter exploration during sparse training. Supticket [34] further enhances the performance of DST by leveraging weight averaging at the late training stage. NeurRev [18] addressed the issue of dormant neurons by removing harmful negative weights. Recent advances include Top-KAST [13], which maintains an auxiliary set of weights for enhanced exploration, and BiDST [14] which formulates DST as a bi-level optimization problem for joint weight-mask optimization. MEST [36] introduces a memory-efficient sparse training framework for edge devices, while Chase [35] adapts dynamic unstructured sparsity into hardware-friendly channel-wise sparsity, both optimizing performance on resource-constrained systems without sacrificing accuracy.
@@
Extreme Sparsity Methods. Most current methods focus on sparsity levels at which pruned networks match or closely approximate the performance of their dense counterparts, a concept referred to as matching sparsity [8]. However, there is growing interest in extreme sparsities, where accuracy trade-offs become inevitable but potentially worthwhile for substantial computational gains. Yet there remains challenges, such as maintaining stable gradient flow and preventing layer collapse, which conventional pruning methods struggle to address. GraSP [33] and SynFlow [31] specifically addressed the gradient preservation and layer collapse issues at high sparsities via careful network initialization. SNIP [17] and SNIP-it [32] proposed PaI methods that showed promising results at higher sparsities, though primarily focused on static masks. De Jorge et al. [2] demonstrated that methods like SNIP and GraSP can perform worse than random pruning at 99% sparsity and beyond, and proposed iterative pruning approaches (Iter-SNIP and FORCE) that maintained meaningful performance up to 99.5% sparsity through gradual parameter elimination. DCTpS [28] introduced ”DCT plus Sparse” layers that combined a fixed dense DCT offset matrix with trainable sparse parameters to maintain propagation at extreme sparsities up to 99.99%.
@@
Activation Functions. Activation functions in sparse networks have traditionally relied on static activation functions such as ReLU [27], which applies a fixed threshold to the input. Parametric variants such as PReLU [11] introduced learnable parameters to adapt the negative slope, offering slight improvements in gradient flow. Swish [29] and Mish [24] aim to smooth gradients for better optimization but remain static throughout training. Dynamic ReLU (DyReLU) [1] introduced an adaptive activation that dynamically adjusts the slopes of a ReLU based on the input.
@@ 3.1 Activation Function
Figure 3: Positive pre-activations analysis in ResNet-34 at 99.99% sparsity. The left figure shows layerwise-comparison of positive pre-activations after DyReLU is completely converted to ReLU. The right figure shows their overall amount before, during and after DyReLU phasing.
In this work, when initializing a network, we replace the standard ReLU activation with a non-ReLU activation. We tested several activations and found that DyReLU worked the best, specifically the DyReLU-B variant, which dynamically adapts and adjusts channel-wise activation coefficients during training, allowing for richer expressivity during early training. As training progresses, we gradually transition from DyReLU to standard ReLU, eliminating the additional parameters. Preliminary details about DST, DyReLU and details on phasing out DyReLU can be found in the Supplementary Materials.
Understanding Gradient Flow Through Neuron Activity. With ReLU, non-positive pre-activations would result in zero gradient and prevent the learning signal with ReLU. To better understand why DyReLU phasing improves performance at extreme sparsity, we analyzed pre-activation values throughout training. As shown in Figure 3, the model trained solely with ReLU maintains around 1.3% positive pre-activations throughout training, with many deeper layers showing none, interrupting gradient propagation.
In contrast, our DyReLU phasing approach naturally begins from a high 56.8%, decreases during the transition phase, and stabilizes at 4.5% even after DyReLU is completely converted to ReLU, three times higher than with ReLU alone. The deeper layers also retain some positive pre-activations, preserving critical activation pathways that would otherwise be lost. This allows extremely sparse networks to continue learning where conventional approaches would struggle due to interrupted gradient propagation. We detail the ablation studies on the effect of DyReLU phasing in section 4.2.
@@ 3.2 Weight Sharing
Much research in sparse training has been about mask optimization - determining the best subset of parameters to keep. However, at extreme sparsities, gradient flow can collapase even with an optimal parameter subset. Therefore, we ask a different question: Can we increase the number of parameter paths in the computational graph while maintaining the same number of learnable parameters?
Our weight sharing mechanism strategically reuses parameters within each layer’s residual blocks, allowing the remaining active parameters to participate in multiple paths of the computational graph. If at a certain sparsity, we have $n$ remaining parameters in the original model, we keep exactly this many unique parameters to learn, but some of these now appear in multiple locations in the computational graph, multiplying their contribution without increasing storage requirements. During backpropagation, shared parameters recieve gradient signals from multiple positions, strengthening gradient flow.
Figure 4: A layer in ResNet-34 with 4 blocks. Block 3 and Block 4 share parameters (and the masks) with block 2. Specifically, conv1 layer (green) of block 3 reuses conv1 layer of block 2, and is multiplied by a learnable scaling factor; similarly, its conv2 layer (blue) reuses conv2 layer of block 2, and is multiplied by another scaling factor.
Specifically, we choose a block inside a layer and share its parameters (along with their masks) with all blocks after it. These later blocks reference the single block tensor that they share with, and multiply by a learnable scaling factor for forward and backward passes. For example, in a ResNet layer with 4 blocks (Figure 4), if the second block shares parameters with blocks 3 and 4, a parameter in block 2 effectively appears three times in the computation graph while only being stored and optimized once.
Thus, we adjust sparsity calculation to account for weight sharing, such that sparsity $\text{sparsity s}=1-\frac{\lVert\mathcal{M}\odot\theta_{\text{learnable}}\rVert_{0}}{\lVert\theta\rVert_{0}}$, where $\lVert$.$\rVert$ . ∥ is the $\ell_{0}$-norm, $\theta_{\text{learnable}}$ is the unique learnable parameter count, and $\theta$ is the theoretical parameter count in the original network with a mask $\mathcal{M}\in\{0,1\}^{|\theta|}$. Formally, let $L$ be the number of residual blocks in a layer, and $R$ be the block that shares parameters ($1<R\leq L$), where the remaining $L-R$ blocks share parameters with the $R$-th block. Let $\theta_{i}$ represent the parameters of the $i$-th block. Our weight sharing scheme can be expressed as:
@@ 3.3 Cyclic Sparsity
While standard DST methods such as SET and RigL adjust the sparsity pattern, sparsity $s$ is consistent throughout training - pruning and regrowing of weights take place at the same $t$. i.e., $\mathcal{M}(t)=\text{Prune}(\theta(t))\cup\text{Grow}(\nabla_{\theta}\mathcal{L})$. In our work, we introduce a cyclic sparsity schedule where the network’s sparsity level itself evolves dynamically during training. Instead of maintaining a constant sparsity, we cyclically adjust the sparsity $s(t)$ between a maximum value $s_{\text{max}}$ and a minimum value $s_{\text{min}}$. The sparsity level $s(t)$ is governed by a cosine schedule:
@@ 4 Experiments
Evaluation Protocol. Our experiments include image classification using ResNet-34 and ResNet-50 [12] on the benchmark datasets CIFAR-10 and CIFAR-100, as well as ResNet-50 on ImageNet, following the same evaluation protocol as other SOTA methods in terms of datasets and backbones. We compare our method with DST methods SET[25] and RigL[4], and a competitive SST method SynFlow[31], which also experiments at very high sparsities. For SynFlow, we use the official repository with its default hyperparameters to evaluate its performance. For SET and RigL, we use the repository provided by ITOP [21]. We keep all hyperparameters as they are optimally configured in the paper, including $\Delta T$, ERK initialization, and the mask update interval of 1500 and 4000 for SET and RigL, respectively.
@@
Method Inference FLOP (M) Inference Network Size
99.95% 99.99% Comp. Cost Theoretical GPU-Supported GPU-Supported Params
ResNet-50 (Dense) 1297.83 - $\mathcal{O}(mn)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ 23.5M
SET 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
RigL 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
EAST (w/o WS) 1.24 (0.001$\times$) 0.26 (0.0002$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
EAST (w/ WS) 2.59 (0.002$\times$) 0.52 (0.0004$\times$) $\mathcal{O}(pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N-N_{s})$ 13.9M
DCTpS+RigL 40.26 (0.031$\times$) 39.09 (0.030$\times$) $\mathcal{O}(q\log q+pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
DCTpS+RigL+EAST 40.26 (0.031$\times$) 39.09 (0.030$\times$) $\mathcal{O}(q\log q+pmn)$ $\mathcal{O}(PN)$ $\mathcal{O}(N)$ 23.5M
Table 2: Comparison of computational complexity and network size of ResNet-50 on different methods. FLOPs are given alongside multiplicative change from dense model ($\times$). Let $N$ denote the total parameters in ResNet-50, $N_{s}$ the total shared parameters, $P\in(0,1)$ the global density, and $m\times n$ the size of flattened weight tensors with density $p$, where $q=\max(m,n)$. ”GPU-Supported Params” refers to the actual number of parameters on a commercial GPU without native support for irregular sparse patterns.
@@
CIFAR-10 and CIFAR-100. We evaluate EAST with ResNet-34 and ResNet-50 backbones, with a particular focus on extreme sparsities (99%-99.99%). We repeat our experiments three times and report the average accuracy in Table 1 (variance visualized in Figure 2), which presents results across different configurations.
At 99%, all dynamic training methods achieve comparable performance, significantly outperforming SynFlow. This is unsurprising as a static mask is expected to lead to worse performance than a dynamic mask [4, 26]. EAST’s advantages become increasingly pronounced at higher sparsities. At 99.90% and 99.95%, EAST beats the next best performing method RigL by a clear margin ranging between 1-2% with ResNet-34 on both datasets, showing significant gains in performance. The difference becomes even more clear with ResNet50, where other methods take a huge plummet on accuracy and EAST shows a more gradual decline. EAST displays an improvement of 38-40% over RigL on CIFAR-10, and 1-7% on CIFAR-100 at 99.90% and 99.95%. The most striking difference appears at 99.99%. We observe that all other methods result in complete collapse to random performance on either dataset and model. In contrast, EAST maintains meaningful accuracies except for ResNet50 on CIFAR-100. For instance, EAST achieves 70.57% on CIFAR-10 compared to 10% random baselines. This pattern holds across both models, though ResNet50 shows more degradation at the highest sparsity levels, presumably due to more prominent vanishing gradient problems in a deeper network, with even fewer parameters spread across each layer.
We evaluate DCTpS and find that while DCTpS demonstrates a clear advantage at 99.95% and 99.99%, it performs worse than SET, RigL and EAST at 99% and sometimes 99.9%. For instance, EAST achieves 3-8% higher accuracy than DCTpS at 99%, without involving additional overhead. To test the versatility of EAST, we further combine it with DCTpS and observe that EAST provides consistent improvements across all configurations tested.
99.50% Sparsity 99.90% Sparsity
Method Top-1 (%) Top-5 (%) Top-1 (%) Top-5 (%)
@@
DCT+RigL 19.99 43.34 6.62 18.84
EAST (ours) 34.45 62.44 11.23 29.47
Inference Performance
Throughput (imgs/sec) Latency (ms)
DCT+RigL 1229.25 10.18
EAST (ours) 1638.41 6.80
@@
Table 3: Performance comparison on ImageNet at extreme sparsity levels using ResNet-50. For throughput and latency measurements, we use batch sizes of 2 and run 100 times and report the average.
@@
ImageNet. Prior methods like SynFlow and DCTpS were evaluated primarily on smaller datasets like CIFAR and Tiny-ImageNet. We extend these baselines to ImageNet by implementing and adapting their published code and evaluate the scalability of EAST on ImageNet using ResNet-50 at sparsity 99.5% and 99.9%. Given the much shortened training epochs for ImageNet, we implement only the DyReLU phasing and weight sharing components, excluding cyclic sparsity. As presented in Table 3, EAST has demonstrated capability in maintaining network robustness on large-scale dataset, outperforming previous SOTA by a significant margin, achieving approximately 2% and 2.7% higher top-1 accuracy than RigL. This advantage is more notable when comparing to other methods including SynFlow and SET, which fail to maintain meaningful performance and collapase to random accuracy.
@@
Interestingly, we observe different dynamics on ImageNet compared to CIFAR experiments. While DCTpS showed competitive advantages on smaller datasets, it does not exhibit the same competitive advantages as seen before, even with the extra overhead. DCTpS alone results in performance collapse and it performs worse than RigL. When combined with RigL, the accuracy improvements are modest and may be attributed primarily to RigL’s robust dynamic sparse exploration policy, rather than the DCT offset. On the inference efficiency front, EAST demonstrates practical advantages. Unlike DCTpS, which relies on a dense DCT matrix, EAST does not have such a drawback, as all DyReLU activations have been replaced with regular ReLU and thus runs just like the original ResNet model with ReLU. Therefore, EAST should theoretically have similar inference speed as the other DST methods, but faster than DCTpS for inference. To verify this, we ran both models for inference 100 times, taking their average and finding that DCTpS has 1.5$\times$ higher latency and 1.33$\times$ lower throughput than ours.
4.1 Complexity Analysis
The operations in neural networks can be framed in terms of matrix multiplication, allowing us to analyze the complexity of these methods. We report the FLOP following RigL’s computation method in Table 2. If we examine sparse matrix with $pmn$ non-zeros, then theoretically the inference cost of EAST without weight sharing and other standard DST methods is just the cost of the sparse matrix, $\mathcal{O}(pmn)$, and their storage requirement is just $\mathcal{O}(PN)$. With weight sharing, EAST maintains the theoretical network size $\mathcal{O}(PN)$ while physically storing $\mathcal{O}(N-N_{s})$ parameters, resulting in built-in compression on GPU’s without native support for irregular sparsity patterns. It reduces the parameter count including zeros from 23.5M to 13.9M, but slightly increases the theoretical inference FLOP count due to the reuse of learnable parameters. This trade-off potentially offers practical benefits in memory-constrained devices.
While DCTpS achieves higher accuracy than standard DST methods at extreme sparsity levels, it comes with substantial computational overhead that limits its practical deployment benefits. DCTpS uses a dense DCT matrix, which theoretically could have a complexity of $\mathcal{O}(qlogq)$. We found that its theoretical inference FLOPs are about 32X higher than the standard DST at 99.95% and 150X higher at 99.99%. This is because the dense DCT computation dominates at extreme sparsities, becoming an increasingly significant bottleneck as sparsity increases. This additional computational overhead directly explains the inference throughput and latency differences observed in our experiments.